The rate of increase of susceptible AIDS victims is proportional to the number of susceptible
persons and number of infected persons. If there are 0
S susceptible persons and 1 infected
person at a time to then i) set up the equation for the spread of the disease ii) solve the resulting
equation iii) give a physical interpretation to the same by plotting the epidemic curve iv) write the
limitations of the model.
1
Expert's answer
2020-03-10T10:45:18-0400

i)

i - number of infected persons at time t

s - number of susceptible persons at time t

k - konstant

t - time

s0 - initial number of susceptible persons

i0 - initial number of infected persons = 1

i+s=s0+1

di/dt = k*i*s

ii)

di/dt = k*i*(s0+1-i)

di/(k*i*s0+k*i-k*i2)=dt

(ln(i)-ln(1-i+s0))/(k+ks0)+c=t (c - constant)

at t=0 i=1

c=ln(s0)/(k+k*s0)

(ln(i)-ln(1-i+s0)+ln(s0))/(k+k*s0)=t

i = exp[k t + k s0 t](s0+1)/(exp[k t + k s0 t]+s0)

iii)

Let s0=1000, k=1 then i(t) would be



iv)

Limitations:

does not take into account mortality, fertility, virus mutations, etc